OpenStudy (rosales0228):

A businessman bought a car dealership that is incurring a loss of $500,000 a year. He decided to strategize in order to turn the business around. In addition to the $500,000 annual loss, his fixed cost for running the dealership on a monthly basis is $5,000. The number of cars sold per week and their probabilities mimic the outcomes of three coins being flipped. The number of cars sold in a week was observed to be the same as the number of tails that appear when three coins are flipped. See the distribution: Number of Tails 0 1 2 3 Probability 1/8 3/8 3/8 1/8

OpenStudy (rosales0228):

@cj49

OpenStudy (rosales0228):

GIVING MEDAL!!!

OpenStudy (cj49):

wat r d options

OpenStudy (rosales0228):

Given that there are 52 weeks in a year, what is the expected revenue per car (rounded to the nearest dollar) that has to be made in order to break even in the first year? $4,308 $7,179 $5,385 $10,769 $3,590

OpenStudy (cj49):

5000*12 = 60 000 for the year, add in the loss we get 560 000 divided by 52 gives a weekly break even amt: 10 770 the expected probability. 0(1/8)+1(3/8)+2(3/8)+3(1/8)=12/8=1.5 he sells 1.5 cars on average per week. 10770/1.5 = 7180

OpenStudy (rosales0228):

thanks!!